Could some one please help me solve this problem I am totally confused? Two angles are complimentary. The sum of the measure of the first angle and one-fourth the second angel is 84.75 degrees. Find the measures of the angle. Smaller __ other__
Answer by nerdybill(7384)

Could some one please help me solve this problem I am totally confused? Two angles are complimentary. The sum of the measure of the first angle and one-fourth the second angel is 84.75 degrees. Find the measures of the angle. Smaller __ other__
.
two angle that are complementary add up to 90 degrees
.
Let x = first angle
then
90-x = second angle
.
from:"The sum of the measure of the first angle and one-fourth the second angel is 84.75 degrees"
x + (90-x)/4 = 84.75
multiplying both sides by 4:
4x + 90-x = 339
3x + 90 = 339
3x = 249
x = 83 degrees (first angle)
.
second angle:
90-x = 90-83 = 7 degrees